10. The data below are (pretend) for the USA. What was the value for the GDP Deflator in 2019?

Base Year 2018 2019

Qx 10 15 20

Qy 5 10 15

Px $2 $4   $5

Py $5 $8 $10

How much inflation (cumulatively) has there been since the base year? __________
Answer #1

GDP deflator in 2019 = nominal GDP/ real GDP * 100

Nominal GDP in 2019 = market value of goods produced in 2019 and measured in current price = 20 * $5 + 15 * $10 = $100 + $150 = $250.

Real GDP in 2019 = market value of goods produced in 2019 and measured in constant price = 20 * $2 + 15 * $5 = $40 + $75 = $115

GDP deflator in 2019 = nominal GDP/ real GDP * 100 = $250/$115 * 100 = 2.17 * 100 = 217.

Inflation in 2019 from base year = % change in CPI from base year.

CPI in base year = 100.

Inflation in 2019 from base year = % change in CPI from base year = (217 - 100)/100 * 100 = 117/100 * 100 = 117%